OpenStudy (anonymous):

Prove that x^2 is always positive. I understand why, but I don't know how to write out.

OpenStudy (mertsj):

Try an indirect proof....assume that it is negative and show that it leads to a contradiction.

OpenStudy (nenadmatematika):

for every x in R y=x^2 stays equal to zero or bigger than zero

myininaya (myininaya):

what do you mean you don't know how to write x^2>0?

myininaya (myininaya):

and yes because one counterexample for all complex would be if x=i i^2=-1<0

OpenStudy (nenadmatematika):

correct....just look out because it could be equal to zero when x=0

myininaya (myininaya):

yep yep

myininaya (myininaya):

so the best thing to say is x^2>=0 for all reals

OpenStudy (zarkon):

show thats its minumm value is 0

myininaya (myininaya):

which nenad already said

OpenStudy (zarkon):

I butchered the crap out of that word....minimum ;)

myininaya (myininaya):

zarkon you don't have to know how to spell

myininaya (myininaya):

you just have to get it a little right

OpenStudy (zarkon):

good...because I can't

OpenStudy (anonymous):

Here is the problem: f'(x) = (x-a)(x-b)^2 0 < a < b "For which values for a and b is f increasing?" And I can't motivate why (x-b)^2 is always positive. Am I over thinking this?

myininaya (myininaya):

f'=2x=0 => x=0 which means x=0 is a critical number f''=2 which means it is concave up

myininaya (myininaya):

so a and b are critical numbers \[f''=(x-b)^2+(x-a) \cdot 2(x-b)=(x-b)[x-b+2(x-a))]=(x-b)[3x-b-2a]\] \[f''(a)=(a-b)[3a-b-2a]=(a-b)[a-b]=(a-b)^2=(b-a)^2>0\] \[f''(b)=0 \text{ inconclusive }\]

myininaya (myininaya):

b>a so b is not a ever so (b-a)^2 is always positive of course we are pretending we have real a and real b

OpenStudy (anonymous):

Thank you very much!

myininaya (myininaya):

but it seems like the question you asking now is different from the one you are asking earlier \[f'(x)=(x-a)(x-b)^2\] we want f'>0 \[(x-a)(x-b)^2>0\] f' is zero when x=a or x=b |-----|---------|---- 0 a b chose number btw 0 and a like a/2 we want the following to be greater than 0 \[f'(\frac{a}{2})=(\frac{-a}{2})(\frac{a}{2}-b)^2 \] we want to choose a number btw a and b now like (a+b)/2 again we want the following to be greater than 0 \[f'(\frac{a+b}{2})=\frac{-a+b}{2}(\frac{a-b}{2})^2=-\frac{a-b}{2}(\frac{a-b}{2})^2=-(\frac{a-b}{2})^3\] now we choose a number after b like b+a we also want this to be bigger than 0 \[f'(b+a)=(b)(a)^2=a^2b\] no now we need to think of a and b such that all of these inequalities are satisfy

myininaya (myininaya):

let me know if i made a mistake somewhere

myininaya (myininaya):

so we have the following three inequalities: \[\frac{-a}{2}(\frac{a}{2}-b)^2>0, -(\frac{a-b}{2})^3>0 , a^2b>0\]

myininaya (myininaya):

"For which values for a and b is f increasing?" wait i'm confused

OpenStudy (anonymous):

I'm sorry I'll rephrase it

myininaya (myininaya):

f'=0 when x=a ,x=b

myininaya (myininaya):

=> f is not increasing at x=a,x=b

OpenStudy (zarkon):

that is not necessarly true

myininaya (myininaya):

f'=(x-a)(x-b)^2?

myininaya (myininaya):

f'=0 when x=a f'=0 when x=b

OpenStudy (anonymous):

I'm sorry it's in Swedish but the picture might help.

myininaya (myininaya):

zarkon i want a counterexample from you

OpenStudy (anonymous):

Thats correct, x=a and x=b gives the zeros for the derivative

myininaya (myininaya):

but we are to find values a and b so that f is increasing

myininaya (myininaya):

but f'=(x-a)(x-b)^2

myininaya (myininaya):

f'=0 when x=a,b

OpenStudy (zarkon):

\[f(x)=x^3\]

myininaya (myininaya):

oh fine

OpenStudy (anonymous):

Yes I know. And you get extra credit for pointing that out in the solution.

myininaya (myininaya):

no zarkon a>0 b>0

myininaya (myininaya):

0<a<b

OpenStudy (zarkon):

def: a function \(f\) is increasing on and interval \(I\) if for any \(x,y\in I\) with \(x<y\) we have that \(f(x)<f(y)\) \[x^3\] satisfies this definition on\(\mathbb{R}\)

myininaya (myininaya):

but a, b cannot equal zero

OpenStudy (zarkon):

x-a can be

OpenStudy (zarkon):

I was just pointing out that just because the derivative is zero does not mean that the function is not increasing.

myininaya (myininaya):

right i get that part you are correct i didn't think about x^3

OpenStudy (asnaseer):

myininaya - as Zarkon says - f'(x)=0 could be a point of inflection. a function can be increasing/decreasing through a point of inflection

myininaya (myininaya):

and other functions like x^3

OpenStudy (anonymous):

The answers says that (x-b)^2 will always be positive. Because of that (x-a) determines the sign of the derivative. x > a and x != a,b Thats is all thats needed

myininaya (myininaya):

but i wonder if i can think of one example where b>a>0

OpenStudy (anonymous):

But the thing I couldn't solve was: Give a good motivation why (x-b)^2 is always positive

OpenStudy (zarkon):

the function is increasing over b the soluituion should just be x>a

OpenStudy (anonymous):

It's not increasing if x=b or x=a

OpenStudy (zarkon):

if you fix a value of x then f is not increasing(it is just a number)...you need to look over an interval....for some open set containig b the function is increasing...even though the derivative is zero at b

OpenStudy (asnaseer):

look at this as an example: http://www.wolframalpha.com/input/?i=y%3D%28x-1%29%28x-4%29%5E2+for+x%3D0+to+5 the function is increasing from x=4 onwards here 4 is equivalent to point x=b

OpenStudy (zarkon):

i should say open interval

OpenStudy (zarkon):

we ar given the derivative...graph the original function

myininaya (myininaya):

http://www.wolframalpha.com/input/?i=find+a+and+b+such+that+-a%5E2%2F2*%28a%2F2-b%29%5E2%3E0%2C+-%28a-b%29%5E3%2F8%3E0%2C+a%5E2b%3E0 i put in those three inequalities wolfram said there is no solution to that system

OpenStudy (zarkon):

this is never true for b>a>0 \[\frac{-a}{2}(\frac{a}{2}-b)^2>0\]

OpenStudy (anonymous):

The answer doesn't require an interval. Even if x will equal a at some point, it's enough to state that at this point the derivative will be = 0

OpenStudy (anonymous):

Thanks for your help Myininaya

OpenStudy (zarkon):

you gave an interval (2 intervals) as an answer... " x > a and x != a,b Thats is all thats needed"

OpenStudy (anonymous):

Sorry sorry, I didn't think realize that. But that's the correct answer.

OpenStudy (zarkon):

it is not

OpenStudy (asnaseer):

myininaya - your original inequalities were correct - they lead to the correct solution of x > a

myininaya (myininaya):

but -a/2<0 since a>0

myininaya (myininaya):

i don't see a solution to the problem

OpenStudy (asnaseer):

you got this for x < a:\[f'(\frac{a}{2})=(\frac{-a}{2})(\frac{a}{2}-b)^2\]which is always negative, so function cannot be increasing over this region. for x>a and x<b you got:\[f'(\frac{a+b}{2})=-(\frac{a-b}{2})^3=-\frac{a-b}{2}(\frac{a-b}{2})^2=\frac{b-a}{2}(\frac{a-b}{2})^2\]which is always positive since b > a. and finally for x > b you got:\[f'(b+a)=a^2b\]which is always positive since 0 < a < b

OpenStudy (asnaseer):

therefore function is increasing for x > a

myininaya (myininaya):

but want f'>0 always we wanted f to be always increasing and that gets messed up on my f'(a/2)

myininaya (myininaya):

oh we wanted to know where f is increasing?

myininaya (myininaya):

or did we want to choose a and b so that f is increasing everywhere?

myininaya (myininaya):

first problem is x>a but second problem has no solution

OpenStudy (zarkon):

it is not possible to pick an a and b such that f is increasing everywhere

OpenStudy (anonymous):

This is an analyzing problem. We should find when f is increasing and when it's zero

OpenStudy (anonymous):

We don't have to do that

OpenStudy (anonymous):

Just state that if x > a, f is increasing. as long as it isn't = a or b

OpenStudy (anonymous):

But since it can't be > a and = a. x > a and x != b is the solution given in the answer sheet.

OpenStudy (asnaseer):

I think your answer sheet is wrong, it should just be x > a as Zarkon stated earlier

OpenStudy (anonymous):

if x=b then f' = 0 which means not increasing

OpenStudy (asnaseer):